How many mL of hydrogen peroxide are needed to make an 11.3% solution by volume of hydrogen peroxide if you want to make 689 mL of solution?


1
Expert's answer
2022-02-19T06:51:05-0500

Volume of H2O2= (11.3÷100)×689

=77.857ml

77.857 ml is equivalent to 30%

Volume of H2O2 solution =(77.857÷30)×100 = 259.52ml